# LoRI: Reducing Cross-Task Interference in Multi-Task Low-Rank Adaptation

### v1

@@ 3.1 Experimental Setup
We conduct a series of experiments to evaluate LoRI’s effectiveness on single-task and multi-task settings, including adapter merging and continual learning. We focus on four capabilities: (i) Natural Language Understanding (NLU): LoRI is trained on the aggregation of eight NLU datasets (Hu et al., 2023), including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SocialIQA (Sap et al., 2019), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and Winogrande (Sakaguchi et al., 2021). We evaluate accuracy on the individual test split for each dataset. (ii) Mathematical Reasoning (Math): LoRI is trained on the GSM8K (Cobbe et al., 2021) training split and evaluated on the GSM8K test split. (iii) Code Generation (Code): LoRI is trained on CodeAlpaca (Chaudhary, 2023) and evaluated using pass@1, pass@5, and pass@10 on HumanEval (Chen et al., 2021). (iv) Safety: LoRI is trained on Saferpaca (Bianchi et al., 2023), which extends Alpaca-Cleaned (Taori et al., 2023) with 2,000 safety instructions. Safety performance is assessed by measuring the refusal rate on harmful queries from HEx-PHI (Qi et al., 2023).
@@
In single-task experiments, we compare LoRI with full fine-tuning (FFT), LoRA (Hu et al., 2021), and DoRA (Liu et al., 2024). In merging experiments, we compare LoRI merging with several LoRA merging methods, including concatenated merging, linear merging (Ilharco et al., 2022), magnitude pruning, TIES-Merging (Yadav et al., 2023), and DARE (Yu et al., 2024). Magnitude pruning, TIES, and DARE are pruning-based approaches that apply sparsification to the $A$ and $B$ matrices before merging, based on a specified density. Magnitude pruning removes low-magnitude parameters; TIES-Merging further merges weights with consistent signs; and DARE performs random pruning followed by rescaling. For fair comparison, all baseline results are reproduced using a consistent experimental setup.
@@
We use Llama-3-8B (Grattafiori et al., 2024) and Mistral-7B (Jiang et al., 2023) as base models. We conduct all experiments on 8 NVIDIA A5000 GPUs. As shown in Figure 1(a), LoRI maintains strong performance even with 90% sparsity in $B$. To explore the impact of sparsity, we provide two variants of LoRI: LoRI-D, which uses dense $B$ matrices, and LoRI-S, which applies 90% sparsity to $B$. Sparsity is implemented by masking the gradients of $B$ during backpropagation. For optimal performance, we use the entire adaptation dataset as the calibration dataset for each task. Ablation results for calibration are presented in Section 3.5. For consistency, we use the same hyperparameters for LoRA and DoRA as for LoRI-D. Uniform merging weights are used for all adapters. Detailed hyperparameter settings are provided in Appendix D.
@@ 3.2 Single-Task Performance
Table 1 presents single-task results on eight NLU benchmarks, while Table 2 reports single-task performance on the math, code, and safety benchmarks. The low-rank dimension is set to $r=32$, with additional results for $r=64$ provided in Appendix E.1. While full fine-tuning (FFT) updates all model parameters, LoRA and DoRA reduce the number of trainable parameters to approximately 1%. LoRI-D further reduces this to about 0.5% by freezing matrices $A$, and LoRI-S pushes this reduction to 0.05% by applying 90% sparsity to matrices $B$, achieving a 95% reduction in trainable parameters compared to LoRA. Despite tuning fewer parameters, LoRI-D and LoRI-S achieve performance comparable to – and even better than – LoRA and DoRA on NLU, math, code, and safety tasks. LoRI-D generally outperforms LoRI-S slightly, likely due to the extremely limited parameter budget in LoRI-S. Remarkably, LoRI-D and LoRI-S consistently outperform FFT, LoRA, and DoRA on code generation tasks. On HumanEval with Llama-3, LoRI-D achieves a pass@10 score of 63.2%, outperforming LoRA by 24.4%. LoRI-S achieves 59.6% pass@10, exceeding LoRA by 17.3%.
@@ 3.3 Adapter Merging
We consider four heterogeneous tasks for LoRA and LoRI merging: NLU, math, code, and safety. This setting is generally more challenging than merging homogeneous adapters, such as multiple NLU adapters. Table 3 presents results for merging LoRAs and LoRIs on these four tasks. For LoRI, we apply concatenated and linear merging to the LoRI-D and LoRI-S variants. Pruning-based methods such as magnitude pruning, TIES, and DARE are not applied to LoRI, since these methods will prune the $A$ matrices as LoRI already sparsifies $B$, resulting in an inconsistent pruning scheme across $A$ and $B$. Additional results, including experiments on merging three adapters and evaluations of pruning-based merging methods on LoRI, are provided in Appendix E.3 and E.4.
As shown in Table 3, directly merging LoRAs results in substantial performance degradation, particularly for code generation and safety alignment. Although pruning-based methods (e.g., DARE, TIES) improve code performance, they often compromise accuracy on other tasks. In contrast, LoRI achieves consistently strong performance across all tasks. Concatenated merging with LoRI-D achieves the best overall performance, closely matching the single-task baseline, which indicates minimal interference between LoRI adapters. For instance, it achieves 62.2% pass@10 on HumanEval and an 86.6 safety score on HEx-PHI. Despite using only 5% of the parameters of LoRA, LoRI-S retains competitive performance. Notably, on code and safety tasks, concatenated merging with LoRI-S outperforms all LoRA merging methods. Linear merging with LoRI also performs competitively, though it lags slightly behind concatenation due to cross-term interactions that introduce some interference. Notably, LoRI eliminates the need for manual selection of merging methods: simple concatenation yields strong results. The choice between LoRI-D and LoRI-S can then be guided by the desired trade-off between performance and parameter efficiency.
@@ 3.4 Continual Learning
Figure 3: Continual learning results from safety to NLU, math, and code domains. Results for NLU are averaged over eight tasks. GSM8K accuracy and HumanEval pass@10 are reported individually. Base model: Llama-3-8B, rank $r=32$.
While merging adapters enables multi-task capabilities, it falls short of providing robust safety alignment in scenarios that demand strong safety guarantees. As shown in Table 3, the highest safety score achieved through LoRA or LoRI merging is 86.6. To address this limitation, we adopt a two-phase training process: first, a safety adapter is trained on the safety alignment dataset Saferpaca; then, it is individually adapted to each downstream task, including NLU, math, and code. Figure 3 presents results from these continual learning experiments. LoRA exhibits severe catastrophic forgetting on safety alignment – particularly in the safety $\rightarrow$ NLU experiment – likely due to the large size of the NLU training split ($\sim$170k examples). Among all methods, LoRI-S achieves the best preservation of safety alignment, even outperforming single-task LoRI-D. This is due to its 90% sparsity in the $B$ matrices, which enables isolated parameter updates between the safety alignment and task adaptation. LoRI-D also shows some resistance to forgetting, benefiting from frozen $A$ matrices. For task adaptation, LoRI-D generally outperforms LoRI-S, as the latter’s aggressive sparsity limits its adaptation capacity. Overall, LoRI offers a lightweight and effective approach to building safety adapters that preserve alignment while supporting downstream task adaptation.
@@ 3.5 Ablation Studies
We use global masks in our experiments that retain the highest-magnitude parameters across all layers and projections. Figure 4(b) presents the sparsity ratios of different projection types (e.g., up, down, key, value) across layers under a 90% sparsity on GSM8K. We observe that feedforward (FFN) projections tend to retain more parameters (i.e., lower sparsity) than self-attention projections, indicating they are more critical for adaptation. Additionally, the top layers are less sparse than lower layers, suggesting that the top layers play a more important role in adaptation.
@@
We adopt uniform weights across all adapters for adapter merging, rather than task-specific weights, as we do not wish to prioritize any individual task. Figure 4(d) shows the effect of different merging weights (0.2, 0.3, 0.4) for concatenated merging with LoRI-S. We observe that LoRI is moderately sensitive to merging weights, with a noticeable trade-off between performance on code tasks and other domains. We adopt 0.3 for all tasks in LoRI-S merging, as it offers a balanced performance across domains.
@@ 4 Conclusion
We identify several promising avenues for extending this work. While LoRI currently leverages unstructured magnitude-based sparsity, future research can explore structured sparsity patterns – such as block sparsity, head pruning, or group-wise masking – which may offer better hardware compatibility. Additionally, although this study focuses on LLMs, the core design of LoRI is modality-agnostic. Extending LoRI to diffusion and vision-language models for multi-modal generation is a promising direction, given the growing impact of adapter-based fine-tuning.
@@ Appendix A Related Works
Parameter-efficient fine-tuning (PEFT) methods for LLMs (Houlsby et al., 2019; Pfeiffer et al., 2020; Li & Liang, 2021; Lester et al., 2021; Liu et al., 2021; Hu et al., 2021) have garnered increasing attention, driving a wide range of algorithmic and architectural advancements. Among them, LoRA (Hu et al., 2021) introduces trainable low-rank matrices into each projection, which can be merged into the pretrained weights. Due to its strong performance and high efficiency, LoRA has become one of the most widely adopted PEFT methods. Several studies have proposed LoRA variants aimed at enhancing learning capacity to resemble FFT (Liu et al., 2024; Nikdan et al., 2024), reducing the number of trainable parameters (Kopiczko et al., 2023; Ding et al., 2023; Zhang et al., 2023b), implementing adaptive parameter budget allocation (Zhang et al., 2023a; d), and integrating LoRA with techniques such as quantization (Dettmers et al., 2024; Xu et al., 2023; Guo et al., 2023) and pruning (Zhang et al., 2023c). Unlike previous methods, LoRI leverages the sparsity of matrices $B$ by applying task-specific masks while keeping matrices $A$ frozen. This significantly reduces the number of trainable parameters while preserving the capabilities of the pretrained model.
@@
Achieving multi-task capabilities typically involves training on a mixture of diverse task datasets (Caruana, 1997; Sener & Koltun, 2018), which is often prohibitively expensive in time and compute. As an alternative, model merging has gained attention for combining multiple task-specific models into a single model (Matena & Raffel, 2022; Ilharco et al., 2022; Yadav et al., 2023; Yu et al., 2024). Fisher Merging (Matena & Raffel, 2022) uses weights from the Fisher information matrix to combine parameters, while Task Arithmetic (Ilharco et al., 2022) employs predefined scaling factors. TIES-Merging (Yadav et al., 2023) prunes low-magnitude parameters and merges those with consistent signs, and DARE (Yu et al., 2024) applies random pruning with rescaling. However, identifying the optimal merging method often requires trial and error. More recently, there has been growing interest in merging task-specific LoRA adapters (Chronopoulou et al., 2023; Huang et al., 2023; Wu et al., 2024; Wang et al., 2024a; Tian et al., 2024; Prabhakar et al., 2024), often utilizing Mixture-of-Experts (MoE) architectures. Nonetheless, these methods typically require additional training to coordinate the adapters effectively. In contrast, LoRI eliminates the need for manual selection of merging methods or additional training. By ensuring approximate orthogonality between adapters, LoRI minimizes interference and preserves task-specific performance.
@@
Catastrophic forgetting is a fundamental challenge in continual learning (McCloskey & Cohen, 1989; Ramasesh et al., 2021; Wang et al., 2024b), where neural networks struggle to retain previously learned knowledge when adapting to new tasks. Wu et al. (2022) analyzed this phenomenon using layer-wise and task-wise probing to assess knowledge retention across tasks. Several studies (Dong et al., 2023; Luo et al., 2023) have empirically examined catastrophic forgetting in the continual fine-tuning of LLMs. To mitigate catastrophic forgetting, various approaches have been proposed. Rehearsal-based methods (Rolnick et al., 2019; Shin et al., 2017) store or generate past data to reinforce prior knowledge during training. Parameter isolation methods (Rusu et al., 2016; Mallya & Lazebnik, 2018; Konishi et al., 2023; Panda et al., 2024) allocate separate subnetworks or sparsely mask parameters for different tasks to prevent interference. Additionally, O-LoRA (Wang et al., 2023) learns tasks in distinct low-rank vector subspaces while ensuring orthogonality between them. LoRI falls under the category of parameter isolation methods, leveraging sparse task-specific masks to mitigate catastrophic forgetting during continual learning.

### v2

@@ 3.1 Experimental Setup
We conduct a series of experiments to evaluate LoRI’s effectiveness on single-task and multi-task settings, including adapter merging and continual learning. We focus on four capabilities: (i) Natural Language Understanding (NLU): LoRI is trained on the aggregation of eight NLU datasets (Hu et al., 2023), including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SocialIQA (Sap et al., 2019), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and Winogrande (Sakaguchi et al., 2021). We evaluate accuracy on the individual test split for each dataset. (ii) Mathematical Reasoning (Math): LoRI is trained on the GSM8K (Cobbe et al., 2021) training split and evaluated on the GSM8K test split. (iii) Code Generation (Code): LoRI is trained on CodeAlpaca (Chaudhary, 2023) and evaluated using pass@1, pass@5, and pass@10 on HumanEval (Chen et al., 2021). (iv) Safety Alignment (Safety): LoRI is trained on Saferpaca (Bianchi et al., 2023), which extends Alpaca-Cleaned (Taori et al., 2023) with 2,000 safety instructions. Safety performance is assessed by measuring the refusal rate on harmful queries from HEx-PHI (Qi et al., 2023).
@@
In single-task experiments, we compare LoRI with full fine-tuning (FFT), LoRA (Hu et al., 2021), and DoRA (Liu et al., 2024). Results for additional PEFT baselines, including VeRA (Kopiczko et al., 2023), IA3 (Liu et al., 2022), LoRA-FA (Zhang et al., 2023b), AdaLoRA (Zhang et al., 2023d), rsLoRA (Kalajdzievski, 2023), PiSSA (Meng et al., 2024), and LoRA+ (Hayou et al., 2024), are available in Appendix E.1. In merging experiments, we compare LoRI merging with several LoRA merging methods, including concatenated merging, linear merging (Ilharco et al., 2022), magnitude pruning, TIES-Merging (Yadav et al., 2023), and DARE (Yu et al., 2024). Magnitude pruning, TIES, and DARE are pruning-based approaches that apply sparsification to the $A$ and $B$ matrices before merging, based on a specified density. Magnitude pruning removes low-magnitude parameters; TIES-Merging further merges weights with consistent signs; and DARE performs random pruning followed by rescaling. For fair comparison, all baseline results are reproduced using a consistent experimental setup.
@@
We use Llama-3-8B (Grattafiori et al., 2024) and Mistral-7B (Jiang et al., 2023) as base models. We conduct all experiments on 8 NVIDIA A5000 GPUs. To explore the impact of sparsity, we provide two variants of LoRI: LoRI-D, which uses dense $B$ matrices, and LoRI-S, which applies 90% sparsity to $B$. Sparsity is implemented by masking the gradients of $B$ during backpropagation. For optimal performance, we use the entire adaptation dataset as the calibration dataset for each task. Ablation results for calibration are presented in Section 3.5. For consistency, we use the same hyperparameters for PEFT baselines as for LoRI-D. For all adapter merging experiments, uniform weights $\alpha_{t}$ are employed across all adapters. The weights $\alpha_{t}$ are treated as hyperparameters, and their ablation study is detailed in Section 3.5. Detailed hyperparameter settings are provided in Appendix D.
@@ 3.2 Single-Task Performance
Table 1 presents single-task performance on eight NLU benchmarks, while Table 2 reports single-task performance on the math, code, and safety benchmarks. Results for additional PEFT baselines are available in Appendix E.1. The rank for our experiments is set to $r=32$. We observed stable performance across different ranks, with additional results for $r=64$ provided in Appendix E.2.
While full fine-tuning (FFT) updates all model parameters, LoRA and DoRA reduce the number of trainable parameters to approximately 1%. LoRI-D further reduces this to about 0.5% by freezing matrices $A$, and LoRI-S pushes this reduction to 0.05% by applying 90% sparsity to matrices $B$, achieving a 95% reduction in trainable parameters compared to LoRA. Despite tuning fewer parameters, LoRI-D and LoRI-S achieve performance comparable to – and even better than – LoRA and DoRA on NLU, math, code, and safety tasks. LoRI-D generally outperforms LoRI-S slightly, due to the extremely limited parameter budget in LoRI-S. Remarkably, LoRI-D and LoRI-S consistently outperform FFT, LoRA, and DoRA on code generation tasks. On HumanEval with Llama-3, LoRI-D achieves a pass@10 score of 63.2%, outperforming LoRA by 24.4%. LoRI-S achieves 59.6% pass@10, exceeding LoRA by 17.3%.
@@ 3.3 Adapter Merging
We consider four heterogeneous tasks for LoRA and LoRI merging: NLU, math, code, and safety. This setting is generally more challenging than merging homogeneous adapters, such as merging multiple NLU adapters. Table 3 presents results for merging LoRAs and LoRIs on these four tasks. For LoRI, we apply concatenated and linear merging to the LoRI-D and LoRI-S variants. Pruning-based methods such as magnitude pruning, TIES, and DARE are not applied to LoRI, since these methods will prune the $A$ matrices as LoRI already sparsifies $B$, resulting in an inconsistent pruning scheme across $A$ and $B$. Additional results, including experiments on merging three adapters and evaluations of pruning-based methods on LoRI, are provided in Appendix E.4 and E.5.
As shown in Table 3, directly merging LoRAs results in substantial performance degradation, particularly for code generation and safety alignment. Although pruning-based methods (e.g., DARE, TIES) improve code performance, they often compromise accuracy on other tasks. In contrast, LoRI achieves consistently strong performance across all tasks.
Concatenated merging with LoRI-D achieves the best overall performance, closely matching the single-task baseline, which indicates minimal interference between LoRI adapters. For instance, it achieves 62.2% pass@10 on HumanEval and an 86.6% refusal rate on HEx-PHI. Despite using only 5% of the parameters of LoRA, LoRI-S retains competitive performance. Notably, on code and safety tasks, concatenated merging with LoRI-S outperforms all LoRA merging methods.
Linear merging with LoRI also performs competitively, though it lags slightly behind concatenated merging due to cross-term interactions that introduce some interference. LoRI eliminates the need for manual selection of merging methods: simple concatenated merging yields strong results. The choice between LoRI-D and LoRI-S can then be guided by the desired trade-off between performance and parameter efficiency. We also note an important trade-off between code generation performance and other domains during adapter merging, a phenomenon further explored in Section 3.5.
@@ 3.4 Continual Learning
Figure 3: Continual learning results from safety to NLU, math, and code domains. Results for NLU are averaged over eight tasks. GSM8K accuracy, HumanEval pass@10, and HEx-PHI refusal rate are reported individually. Base model: Llama-3-8B, rank $r=32$.
While merging adapters enables multi-task capabilities, it falls short of providing robust safety alignment in scenarios that demand strong safety guarantees. As shown in Table 3, the highest refusal rate on HEx-PHI achieved through LoRA or LoRI merging is 86.6%. To address this limitation, we adopt a two-phase training process: first, a safety adapter is trained on the safety alignment dataset Saferpaca; then, it is individually adapted to each downstream task, including NLU, math, and code.
Figure 3 presents results from these continual learning experiments. LoRA exhibits severe catastrophic forgetting on safety alignment – particularly in the safety $\rightarrow$ NLU experiment – likely due to the large size of the NLU training split ($\sim$170k examples). Among all methods, LoRI-S achieves the best preservation of safety alignment, even outperforming single-task LoRI-D. This is due to its 90% sparsity in the $B$ matrices, which enables isolated parameter updates between the initial safety alignment and subsequent task adaptations. LoRI-D also shows some resistance to forgetting, benefiting from frozen $A$ matrices. For task adaptation, LoRI-D generally outperforms LoRI-S, as the latter’s aggressive sparsity limits its adaptation capacity. Overall, LoRI offers a lightweight and effective approach to building safety adapters that preserve alignment while supporting adaptation to downstream tasks.
@@ 3.5 Ablation Studies
We use model-wise masks in our experiments that retain the highest-magnitude parameters across all layers and projections. Figure 4(b) presents the sparsity ratios of different projection types (e.g., up, down, key, value) across layers under a 90% sparsity on GSM8K. We observe that feedforward (FFN) projections tend to retain more parameters (i.e., lower sparsity) than self-attention projections, indicating they are more critical for adaptation. Additionally, the top layers are less sparse than lower layers, suggesting that the top layers play a more important role in adaptation.
@@
We adopt uniform weights across all adapters for adapter merging, rather than task-specific weights, as we do not wish to prioritize any individual task. Figure 4(d) shows the effect of different merging weights (0.2, 0.3, 0.4) for concatenated merging with LoRI-S. We observe that LoRI is moderately sensitive to merging weights, with a noticeable trade-off between performance on code tasks and other domains. We adopt 0.3 for all adapters in LoRI-S merging, as it offers a balanced performance across domains.
@@ 4 Conclusion
We identify several promising avenues for extending this work. While LoRI currently leverages unstructured magnitude-based sparsity, future research can explore structured sparsity patterns – such as block sparsity, head pruning, or group-wise masking – which may offer better hardware compatibility. Additionally, although this study focuses on LLMs, the core design of LoRI is modality-agnostic. Extending LoRI to diffusion and vision-language models for multi-modal generation is a promising direction.
@@ Appendix A Related Works
Parameter-efficient fine-tuning (PEFT) methods for LLMs (Houlsby et al., 2019; Pfeiffer et al., 2020; Li & Liang, 2021; Lester et al., 2021; Liu et al., 2021; Hu et al., 2021) have received increasing attention in recent years. Among them, LoRA (Hu et al., 2021), which introduces trainable low-rank matrices, has become one of the most widely adopted PEFT methods due to its strong performance and efficiency. LoRI is motivated by reducing parameter redundancy in LoRA through an asymmetric design: we freeze the projection matrices $A$ and enforce sparsity on the matrices $B$. Our work is closely related to several lines of research. In terms of parameter efficiency, our goal is shared by methods such as IA3 (Liu et al., 2022), VeRA (Kopiczko et al., 2023), and FourierFT (Gao et al., 2024). More specifically, our approach builds on the concept of asymmetric LoRA variants, which has been explored in works like LoRA-FA (Zhang et al., 2023b), AsymmetryLoRA (Zhu et al., 2024), and HydraLoRA (Tian et al., 2024). However, LoRI is distinct from these works by uniquely combining frozen $A$ with sparsely updated $B$. This targeted, asymmetric pruning of only the $B$ matrices also differentiates our method from general LoRA pruning techniques like Loraprune (Zhang et al., 2023c), LoRA-drop (Zhou et al., 2024), and SoRA (Ding et al., 2023), as well as SVD-based approaches such as AdaLoRA (Zhang et al., 2023d) and PiSSA (Meng et al., 2024).
@@
Achieving multi-task capabilities typically involves training on a mixture of diverse task datasets (Caruana, 1997; Sener & Koltun, 2018), which is often prohibitively expensive in time and compute. As an alternative, model merging has gained attention for combining multiple task-specific models into a single model (Matena & Raffel, 2022; Ilharco et al., 2022; Yadav et al., 2023; Yu et al., 2024). Fisher Merging (Matena & Raffel, 2022) uses weights from the Fisher information matrix to combine parameters, while Task Arithmetic (Ilharco et al., 2022) employs predefined scaling factors. TIES-Merging (Yadav et al., 2023) prunes low-magnitude parameters and merges those with consistent signs, and DARE (Yu et al., 2024) applies random pruning with rescaling. However, identifying the optimal merging method often requires trial and error. More recently, there has been growing interest in merging task-specific LoRA adapters (Chronopoulou et al., 2023; Huang et al., 2023; Wu et al., 2024; Wang et al., 2024a; Prabhakar et al., 2024; Stoica et al., 2024), often utilizing Mixture-of-Experts (MoE) architectures. Nonetheless, these methods typically require additional training to coordinate the adapters effectively. In contrast, LoRI eliminates the need for manual selection of merging methods or additional training. By ensuring approximate orthogonality between adapters, LoRI minimizes interference and preserves task-specific performance.
@@
Catastrophic forgetting is a fundamental challenge in continual learning (McCloskey & Cohen, 1989; Ramasesh et al., 2021; Liang et al., 2023; Wang et al., 2024b), where neural networks struggle to retain previously learned knowledge when adapting to new tasks. Wu et al. (2022) analyzed this phenomenon using layer-wise and task-wise probing to assess knowledge retention across tasks. Several studies (Dong et al., 2023; Luo et al., 2023) have empirically examined catastrophic forgetting in the continual fine-tuning of LLMs. To mitigate catastrophic forgetting, various approaches have been proposed. Rehearsal-based methods (Rolnick et al., 2019; Shin et al., 2017) store or generate past data to reinforce prior knowledge during training. Parameter isolation methods (Rusu et al., 2016; Mallya & Lazebnik, 2018; Konishi et al., 2023; Panda et al., 2024) allocate separate subnetworks or sparsely mask parameters for different tasks to prevent interference. Additionally, O-LoRA (Wang et al., 2023) learns tasks in distinct low-rank subspaces while ensuring orthogonality between them. LoRI falls under the category of parameter isolation methods, leveraging sparse task-specific masks to mitigate catastrophic forgetting during continual learning.